OpenStudy (fellowroot):

Find a sequence for the following terms. 2,6,12,20,30,42,56,72,90,110,132

OpenStudy (fellowroot):

when i tried to do it i found out that the differencces are 4,6,8,10,12 and the differences of those differences are 2, 2, 2, 2, 2, I know this means something, but I simply forgot how to come up with the sequence. its doesnt really have a common difference yet i dont think its geometric either.

OpenStudy (mathmath333):

u mean u want to find the \(n^{th}\)term

OpenStudy (fellowroot):

yea a general formula

ganeshie8 (ganeshie8):

you have almost worked it

ganeshie8 (ganeshie8):

Notice that you get a constant when you differentiate a second degree polynomial

ganeshie8 (ganeshie8):

so look for a quadratic function

OpenStudy (fellowroot):

yes i see a constant from my experience that means its an exponential formula like y=a^x where a is the constant

ganeshie8 (ganeshie8):

\(\large a_n = xn^2 + yn + z\)

OpenStudy (fellowroot):

but also we are taking the difference not differentiating.

ganeshie8 (ganeshie8):

use the terms to find the coefficients x,y,z

OpenStudy (fellowroot):

so make a system of equatoins and solve it in a matrix

ganeshie8 (ganeshie8):

you're right, we are taking differences and not differeintiating. whats the difference between `nth defference` and a `nth derivative` ?

ganeshie8 (ganeshie8):

one is discrete, and the other is continuous. thats all right ?

ganeshie8 (ganeshie8):

yup pick any 3 terms since u have 3 unknonws

OpenStudy (fellowroot):

okay im going to try it.

ganeshie8 (ganeshie8):

you want the sequence to start at n=0 or 1 ?

OpenStudy (fellowroot):

well actually the sequence could start at 2 or even smaller than that 0 (even though i didnt list it above) and im trying to see where i want to start from

ganeshie8 (ganeshie8):

if you have the choice, make it start at 0 then you will get one coeffecient freely : z = 2

OpenStudy (fellowroot):

well i get An=xn^2+yn+z if it starts at zero the original sequence it then becomes 0=x(0)^2+y(0)+z so z=0

ganeshie8 (ganeshie8):

is your first term 0 ?

ganeshie8 (ganeshie8):

\(\large a_0 = 2\) right ?

OpenStudy (fellowroot):

i just figured out that it could be zero the first term

OpenStudy (fellowroot):

which i didnt wright in my original

ganeshie8 (ganeshie8):

hmm the sequence as it stands, the first term is 2

OpenStudy (fellowroot):

well it could start at 2 or start at zero and im trying to figure out which one i want to choose

OpenStudy (fellowroot):

but we can start at Ao=2 as in the original

ganeshie8 (ganeshie8):

okay finding the coefficients should be easy once you figure out that its a quadratic

OpenStudy (fellowroot):

but how did you know it was a quadratic and not for example a cubic function?

ganeshie8 (ganeshie8):

consider a linear function, the slope is constant so the first differences AND the first derivative is constant, right ?

OpenStudy (fellowroot):

yes

OpenStudy (fellowroot):

okay i solved the matrix and it looks as if its an infinitely many solution one

OpenStudy (fellowroot):

oh nevermind let me redo it

ganeshie8 (ganeshie8):

here is the general rule : \(\large n\)th difference of a polynomial of \(\large n\)th degree is constant

OpenStudy (fellowroot):

omg that is so true

OpenStudy (fellowroot):

so the number of times you take differences until you reach a constant tell you what degree the polynomial needs to be

ganeshie8 (ganeshie8):

for a cubic, you will get the 3rd differences constant

ganeshie8 (ganeshie8):

only IF the sequence is can be represented using a polynomial

OpenStudy (fellowroot):

makes sense. but i swear i did a problem where i was taking differences of the differences and it gave the constant value which is also the base value of an exponential function y=a^x. so how would one know if by taking differences of the differences you would get a polynomical or an exponential?

ganeshie8 (ganeshie8):

for an exponential function, the differences are never constant... thats the reason exponential functions can be infinitely differentiable

OpenStudy (fellowroot):

the sequence i get is An=n^2+3n+2 and it looks as if this is the correct sequence!!!

ganeshie8 (ganeshie8):

so using an exponential function to model a polynomial data may not be a good idea

OpenStudy (fellowroot):

thank u for your help since now i got the answer and know i also know a new technique for sequences!!!

OpenStudy (mathmath333):

there is a ready made formula also

OpenStudy (mathmath333):

formual is \(\large\tt \begin{align} \color{blue}{n^{th}term\\ =a+(n-1)d+\dfrac{(n-1)(n-2)c}{2}}\end{align}\) where a = first term d=difference between first two terms c=second difference

OpenStudy (fellowroot):

wow, ive never seen that formula before. its not in my precal book.

ganeshie8 (ganeshie8):

thats a neat formula

ganeshie8 (ganeshie8):

btw, we can work this particular sequence without using differences / formulas

ganeshie8 (ganeshie8):

there is a nice pattern incase if you haven't noticed already

OpenStudy (mathmath333):

but it only works in case the second differences are constant if it exceesd the second difference,then it doesnt work

ganeshie8 (ganeshie8):

2,6,12,20,30,42,56,72,90,110,132 1*2 = 2 2*3 = 6 3*4 = 12 4*5 = 20 ...

OpenStudy (mathmath333):

lol,thats easy

ganeshie8 (ganeshie8):

:)

OpenStudy (fellowroot):

i would assume that if the sequence had to be subtracted 3 times to find a constant then the formula would be An= a+ (n-1)d +(1/2)(n-1)(n-2)c + (1/3)(n-1)(n-2)(n-3)b , where b is the third difference

OpenStudy (mathmath333):

that would be an invention then

ganeshie8 (ganeshie8):

ahh that looks nasty, i would rather use a matrix

OpenStudy (fellowroot):

thanks again everyone